Show that the points (2,0,1), (0,4,-3), (-2,5,0) are non-collinear. Hence the equation of plane passing through them.
1
Expert's answer
2020-11-17T16:25:26-0500

"\textsf{Let}\,\, A = (2, 0, 1), B = (0, 4, 3), C = (-2, 5, 0)\\

\textsf{If}\, A.(BC) = 0,\, \textsf{the three points are collinear.} \\

\begin{aligned}
\begin{vmatrix}
2 & 0 & 1 \\
0 & 4 & 3 \\
-2 & 5 & 0
\end{vmatrix} &= 
2
\begin{vmatrix}
4 & 3 \\
5 & 0
\end{vmatrix} + 
\begin{vmatrix}
0 & 4 \\
-2 & 5
\end{vmatrix}
\\&= 2(-15) + 8 = -30 + 8 = -22
\end{aligned} \\

\textsf{Since}\,\, A.(B\times C) \neq 0, \textsf{the three points are}\\
\textsf{not collinear.}\\

\textsf{Also, since we are not given a normal}\\
\textsf{vector, we need to find one by taking}\\
\textsf{the cross product of the displacement vector}\\
\overrightarrow{u}\,\textsf{from}\,\, A \,\textsf{to}\, B \,\, \textsf{and the}\\
\textsf{displacement vector}\,\,\overrightarrow{v} \textsf{from}\,\,B \, \textsf{to} C. \\

\textsf{Then we obtain a vector}\,\overrightarrow{n} \,\textsf{which is}\\
\textsf{normal (orthogonal) to each of the original}\\
\textsf{vectors (and thus orthogonal to the plane).}\\

AB = (0, 4, 3) - (2, 0, 1) = (-2, 4, 2 ) \\

BC = (-2, 5, 0) - (0, 4, 3) = (-2, 1, -3)\\

\overrightarrow{u} \times \overrightarrow{v} = (-2, 4, 2 ) \times (-2, 1, -3)\\

\begin{aligned}
\overrightarrow{u} \times \overrightarrow{v} &= \begin{vmatrix}
\textbf{i} & \textbf{j} & \textbf{k}\\
-2 & 4 & 2\\
-2 & 1 & -3
\end{vmatrix}
\\&= \textbf{i}(-12 - 2) - \textbf{j}(6 + 4) + \textbf{k}(-2 + 8)
\\&= -14\textbf{i} - 10\textbf{j} + 6\textbf{k}
\end{aligned}\\

\overrightarrow{n} = (-14, -10, 6),\,\, a = (2, 0, 1)\\

\textsf{Hence the equation of the plane is}\\

(\overrightarrow{r} - a)\overrightarrow{n} = 0\\

-14(x - 2) - 10(y - 0) + 6(z - 1) = 0\\

-14x - 10y + 6z = -22"